Circular No. 7455

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 S4 (LINEAR) D. Schleicher, Lowell Observatory; and C. Eberhardy, University of

Washington, used the Hall 1.1-m telescope at Lowell Observatory to obtain narrowband photometry of comet C/1999 S4, with the following averaged results: June 10–12, log Q(OH) = 28.41, log Q(CN) = 25.54, log Afρ = 2.83 (cf. IAUC 7342); July 13 (seven sets), log Q(OH) = 28.24, log Q(CN) = 25.45; log Afρ = 2.55. The equivalent log Q(water; vectorial) is 28.42 for July 13, and no significant temporal or aperture variations were observed. Significant variability was observed during the June observations, with the gas-production rates on June 11 being 30–50 percent larger than on June 10, and then on June 12 dropping to 10 percent less than on June 10.

Additional visual m_{1}, coma-diameter, and tail-length estimates: July 8.73 UT, 7.3, 8′, – (Y. Nagai, Yamanashi, Japan, 7×35 binoculars); 10.11, 7.5, 5′ × 8′, 40′ (K. Sárneczky, Calar Alto, Spain, 8×56 binoculars); 11.12, 7.8, 3′, 1^{o}.2 (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars); 12.99, 7.0, 6′, 0^{o}.6 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 14.97, 6.8, 3′, – (H. Luethen, Altona, Germany, 10×50 binoculars); 15.34, 7.0, 5′.5, – (C. E. Spratt, Victoria, BC, 0.08-m refractor; strong moonlight).

The following ephemeris, computed from new orbital elements including nongravitational parameters as published on MPEC 2000-O02, shows the comet on July 25 to be 0′.7 north of the position given on IAUC 7449:

2000 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 15 4^{h}41.^{m}23 +61^{o}43.^{′}9 0.504 0.796 50.^{o}3 100.^{o}5 6.5
17 5 49.09 +64 34.6 0.450 0.786 47.5 107.5 6.2
19 7 24.51 +64 46.9 0.408 0.778 44.2 114.3 6.0
21 9 02.69 +60 18.3 0.381 0.772 41.2 119.8 5.8
23 10 14.99 +51 37.3 0.373 0.768 39.8 122.1 5.7
25 11 00.96 +41 09.3 0.385 0.765 40.6 120.3 5.8
27 11 30.09 +31 03.0 0.416 0.765 42.9 115.4 5.9
29 11 49.30 +22 22.2 0.461 0.767 45.6 108.9 6.2
31 12 02.52 +15 19.3 0.517 0.771 47.9 102.2 6.4
Aug. 2 12 11.96 + 9 42.6 0.580 0.777 49.6  95.7 6.7
4 12 18.89 + 5 14.7 0.648 0.784 50.6  89.7 7.0
6 12 24.09 + 1 39.7 0.718 0.794 51.1  84.1 7.3
8 12 28.08 − 1 15.0 0.790 0.805 51.2  78.9 7.5
10 12 31.18 − 3 39.1 0.863 0.818 50.9  74.1 7.8
12 12 33.63 − 5 39.5 0.936 0.833 50.4  69.6 8.1
14 12 35.58 − 7 21.4 1.009 0.848 49.6  65.5 8.3
2000 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green